# Fluctuation-induced giant magnetoresistance in charge-neutral graphene

### v1

@@ Abstract
It has been appreciated since the pioneering work of Gurzhi Gurzhi (1968) that charge transport in high-mobility conductors at finite temperature involves the hydrodynamic flow of the electron liquid. In the hydrodynamic regime, the resistance of the system is determined by the viscosity and thermal conductivity of the electron liquid (see, e.g., the reviews Spivak et al. (2010); Narozhny et al. (2017); Lucas and Fong (2018); Levchenko and Schmalian (2020); Narozhny (2022); Fritz and Scaffidi (2024), as well as Refs. Andreev et al. (2011); Levchenko et al. (2017)).
@@
The physical mechanism underlying the fluctuation-driven enhancement of the intrinsic conductivity can be understood as follows. Consider a graphene Hall-bar device shown in Fig. 1. The Johnson–Nyquist noise produces fluctuations of the electron density $\delta n$. The force exerted by an applied electric field $\bm{E}$ on the fluctuating charge density induces a modulation of the hydrodynamic velocity $\delta\bm{u}$ whose magnitude is determined by the balance between viscous force and the driving force. For the Fourier amplitude with a wavevector $q$, this yields $\eta q^{2}\delta\bm{u}_{q}\sim e\bm{E}\delta n_{q}$, where $\eta$ is the shear viscosity. The correlation between the velocity and charge density fluctuations, creates an average advection current proportional to $\bm{E}$, which enhances the conductivity. For Coulomb interaction screened by a gate, charge density fluctuations decay at a rate $\gamma_{q}\propto q^{2}$. The average advection current caused by the fluctuations can then be estimated as $\langle\delta\bm{j}\rangle\simeq e^{2}\bm{E}\sum_{q}\langle\delta n^{2}_{q}\rangle/(\eta q^{2}\gamma_{q})$. The thermal average of the density fluctuations follows from the fluctuation–dissipation theorem, $\langle\delta n^{2}_{q}\rangle\propto T\sigma_{0}q^{2}$, where $\sigma_{0}$ is the intrinsic conductivity of the pristine electron liquid. It follows from these estimates that the fluctuation-induced conductivity diverges logarithmically at long distances. Below we develop a quantitative theory of the fluctuation contribution, $\sigma_{\mathrm{fl}}$, to the system conductivity and its dependence on the magnetic field. We find that the magnetoresistance of the system is dominated by the fluctuation conductivity in a wide interval of the magnetic fields.
The influence of hydrodynamic fluctuations on charge transport can be studied by applying the method developed by Landau and Lifshitz for Newtonian liquids Landau and Lifshitz (1957); Lifshitz and Pitaevskii (2014); Forster (2019) to the case of charge-neutral electron liquids with a nonvanihshing intrinsic conductivity. To this end, Langevin sources are introduced into the hydrodynamic equations Levchenko et al. (2022). We find that the dominant fluctuation corrections to the electrical conductivity arise from the Johnson-Nyquist noise associated with the intrinsic conductivity of the electron liquid. Therefore, below we focus on this contribution and defer the comprehensive treatment of hydrodynamic fluctuations in charge-neutral electron liquids to a separate work.
@@
$\langle\delta j_{\alpha}(\bm{r},t)\delta j_{\beta}(\bm{r}^{\prime},t^{\prime})\rangle=2T\sigma_{0}\delta_{\alpha\beta}\delta(\bm{r}-\bm{r}^{\prime})\delta(t-t^{\prime}).$ (3)
@@
$en_{\bm{q},\omega}\left(\bm{E}-\frac{\bm{q}(\bm{q}\cdot\bm{E})}{q^{2}}\right)-\frac{\sigma_{0}H^{2}}{c}\bm{u}_{\bm{q},\omega}-\eta q^{2}\bm{u}_{\bm{q},\omega}$
@@
Here we introduced the following characteristic energies in the problem: $\gamma_{\sigma}=\frac{\sigma_{0}}{e^{2}}(U(q)+\partial_{n}\mu)q^{2}$ describes attenuation of fluctuations due to the Maxwell mechanism of charge relaxation, similarly $\gamma_{\nu}=\nu q^{2}$, with $\nu=\eta/\rho$ being kinematic viscosity, describes relaxation via viscous decay of vorticity. Finally, $\gamma_{H}=\sigma_{0}H^{2}/c^{2}\rho$ describes attenuation due to magnetic friction Müller and Sachdev (2008).
The fluctuation-induced contribution to the DC current density is then given by
$\langle\delta\bm{j}\rangle=\int\frac{d^{2}qd\omega}{(2\pi)^{3}}\langle en_{-\bm{q},-\omega}\bm{u}_{\bm{q},\omega}\rangle=\sigma_{\text{fl}}\bm{E},$ (9)
@@
An interesting aspect of the main result, Eq. (Fluctuation-induced giant magnetoresistance in charge-neutral graphene), is that in the ideal liquid approximation, where all the dissipative coefficients of the electron liquid are set to zero, the fluctuation contribution to the conductivity diverges as $\sigma_{\text{fl}}\propto 1/\sigma_{0}$. This may seem counterintuitive, since the strength of the Langevin sources vanishes at $\sigma_{0}\to 0$. The reason for this divergence may be understood as follows. The variance of equal-time fluctuations of the charge density, being a thermodynamic quantity, is independent of $\sigma_{0}$. On the other hand, the relaxation time of the charge density fluctuations diverges at $\sigma_{0}\to 0$, resulting in divergent advection current. A similar counterintuitive dependence on the dissipative coefficients of the electron liquid occurs in the resistivity in the presence of long-range disorder Andreev et al. (2011) and in fluctuation-driven thermal drag Levchenko et al. (2022).

### v2

@@ Abstract
It has been appreciated since the pioneering work of Gurzhi Gurzhi (1968) that charge transport in high-mobility conductors at finite temperature involves the hydrodynamic flow of the electron liquid. The hydrodynamic regime is realized in the range of temperatures and sample purities, in which the momentum relaxation length caused by electron-impurity and electron-phonon scattering exceeds the relaxation length due to momentum conserving electron-electron collisions. In this case, the resistance of the system is determined by the viscosity and thermal conductivity of the electron liquid (see, e.g., the reviews Spivak et al. (2010); Narozhny et al. (2017); Lucas and Fong (2018); Levchenko and Schmalian (2020); Narozhny (2022); Fritz and Scaffidi (2024), as well as Refs. Andreev et al. (2011); Levchenko et al. (2017)).
@@
The physical mechanism underlying the fluctuation-driven enhancement of the intrinsic conductivity can be understood as follows. Consider a graphene Hall-bar device shown in Fig. 1. The Johnson–Nyquist noise produces fluctuations of the electron density $\delta n$. The force exerted by an applied electric field $\bm{E}$ on the fluctuating charge density induces a modulation of the hydrodynamic velocity $\delta\bm{u}$ whose magnitude is determined by the balance between viscous force and the driving force. For the Fourier amplitude with a wavevector $q$, this yields $\eta q^{2}\delta\bm{u}_{q}\sim e\bm{E}\delta n_{q}$, where $\eta$ is the shear viscosity. The correlation between the velocity and charge density fluctuations, creates an average advection current proportional to $\bm{E}$, which enhances the conductivity. For Coulomb interaction screened by a gate, charge density fluctuations decay at a rate $\gamma_{q}\propto q^{2}$. The average advection current caused by the fluctuations can then be estimated as $\bm{j}_{\text{fl}}\simeq e^{2}\bm{E}\sum_{q}\langle\delta n^{2}_{q}\rangle/(\eta q^{2}\gamma_{q})$. The thermal average of the density fluctuations follows from the fluctuation–dissipation theorem, $\langle\delta n^{2}_{q}\rangle\propto T\sigma_{0}q^{2}$, where $\sigma_{0}$ is the intrinsic conductivity of the pristine electron liquid. It follows from these estimates that the fluctuation-induced conductivity diverges logarithmically at long distances. Below we develop a quantitative theory of the fluctuation contribution, $\sigma_{\mathrm{fl}}$, to the system conductivity and its dependence on the magnetic field. We find that the magnetoresistance of the system is dominated by the fluctuation conductivity in a wide interval of the magnetic fields.
The influence of hydrodynamic fluctuations on charge transport can be studied by applying the method developed by Landau and Lifshitz for Newtonian liquids Landau and Lifshitz (1957); Lifshitz and Pitaevskii (2014); Forster (2019) to the case of charge-neutral electron liquids with a nonvanihshing intrinsic conductivity. To this end, Langevin sources are introduced into the hydrodynamic equations Levchenko et al. (2022a). We find that the dominant fluctuation corrections to the electrical conductivity arise from the Johnson-Nyquist noise associated with the intrinsic conductivity of the electron liquid. Therefore, below we focus on this contribution and defer the comprehensive treatment of hydrodynamic fluctuations in charge-neutral electron liquids to a separate work.
@@
$\langle\delta\bm{j}_{\alpha}(\bm{r},t)\delta\bm{j}_{\beta}(\bm{r}^{\prime},t^{\prime})\rangle=2T\sigma_{0}\delta_{\alpha\beta}\delta(\bm{r}-\bm{r}^{\prime})\delta(t-t^{\prime}).$ (3)
@@
$en_{\bm{q},\omega}\left(\bm{E}-\frac{\bm{q}(\bm{q}\cdot\bm{E})}{q^{2}}\right)-\frac{\sigma_{0}H^{2}}{c^{2}}\bm{u}_{\bm{q},\omega}-\eta q^{2}\bm{u}_{\bm{q},\omega}$
@@
where
$\gamma_{\sigma}=\frac{\sigma_{0}}{e^{2}}(U(q)+\partial_{n}\mu)q^{2},\quad\gamma_{\nu}=\nu q^{2},\quad\gamma_{H}=\sigma_{0}H^{2}/c^{2}\rho.$ (9)
Here we introduced the following characteristic energies in the problem: $\gamma_{\sigma}$ describes attenuation of fluctuations due to the Maxwell mechanism of charge relaxation, similarly $\gamma_{\nu}$, with $\nu=\eta/\rho$ being kinematic viscosity, describes relaxation via viscous decay of vorticity. Finally, $\gamma_{H}$ describes attenuation due to magnetic friction Müller and Sachdev (2008). The fluctuation-induced contribution to the DC current density is then given by
$\bm{j}_{\text{fl}}=\int\frac{d^{2}qd\omega}{(2\pi)^{3}}\langle en_{-\bm{q},-\omega}\bm{u}_{\bm{q},\omega}\rangle=\sigma_{\text{fl}}\bm{E},$ (10)
@@
An interesting aspect of the main result, Eq. (Fluctuation-induced giant magnetoresistance in charge-neutral graphene), is that in the ideal liquid approximation, where all the dissipative coefficients of the electron liquid are set to zero, the fluctuation contribution to the conductivity diverges as $\sigma_{\text{fl}}\propto 1/\sigma_{0}$. This may seem counterintuitive, since the strength of the Langevin sources vanishes at $\sigma_{0}\to 0$. The reason for this divergence may be understood as follows. The variance of equal-time fluctuations of the charge density, being a thermodynamic quantity, is independent of $\sigma_{0}$. On the other hand, the relaxation time of the charge density fluctuations diverges at $\sigma_{0}\to 0$, resulting in divergent advection current. A similar counterintuitive dependence on the dissipative coefficients of the electron liquid occurs in the resistivity in the presence of long-range disorder Andreev et al. (2011) and in fluctuation-driven thermal drag Levchenko et al. (2022a).
The effect of weak momentum relaxation induced by disorder and electron-phonon scattering can be incorporated into Eq. (Fluctuation-induced giant magnetoresistance in charge-neutral graphene) by introducing a friction force proportional to the fluid velocity, $k\bm{u}_{\bm{q},\omega}$  [^2] For example, in the model of charge puddle disorder Lucas et al. (2016); Li et al. (2020), it can be shown that $k=\frac{e^{2}}{2\sigma_{0}}\langle\delta n^{2}\rangle$, where $\delta n$ denotes density variations induced by the disorder potential and the angular brackets indicate an average over the system.. To leading order, momentum relaxation be accounted for by making the following replacement $\gamma_{H}\to\gamma_{H}+k/\rho$ in our expressions.
Finally, let us compare our mechanism of magnetoresistance (MR) to the previously discussed mechanisms of MR in graphene devices at charge neutrality. If a hydrodynamic flow perpendicular to the direction of the applied electric field is allowed by the sample geometry, as is the case for Corbino samples, the EMF in the frame of the liquid is reduced by the flow, leading to positive MR Müller et al. (2008); Levchenko et al. (2022b); Gall et al. (2023); Levchenko et al. (2024). In our Hall bar geometry, this mechanism is precluded by the boundary conditions. In the absence of a macroscopic flow, bulk Narozhny et al. (2015) and boundary Alekseev et al. (2015) scattering contributions to the MR at charge neutrality were discussed. The characteristic field strengths associated with these mechanisms are independent of the sample size. Our mechanism, produces strong MR at much weaker fields $H_{T}$, which are inversely proportional to the sample size, see in Eq. (15). It also has a markedly different dependence on $H$. This testing enables identifying the contributions of the different MR mechanisms and testing theoretical predictions in experimental.